# Invariant Graph Transformer for Out-of-Distribution Generalization

### v1

@@ Abstract.
copyright: acmlicenseddoi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/2018/06
@@ 1. Introduction
To address these challenges, we propose GOODFormer, Graph Out-Of-Distribution Generalized Transformer [^1] We will release the source code at publication time.. Our proposed method handles the distribution shifts of graphs by capturing invariant subgraphs and learning OOD generalized graph representations while resolving the aforementioned two challenges. Specifically, as shown in Figure 1, our proposed method achieves this goal by jointly optimizing three mutually reinforcing modules. First, we design a GT-based entropy-guided invariant subgraph disentangler to separate invariant and variant subgraphs using attention coefficients. To maintain the sharpness of the attention function, we propose entropy-guided sharp attention, comprising an entropy-guided regularizer and a lightweight temperature-based test-time training method. Next, we introduce an evolving subgraph positional and structural encoder to learn PSEs of dynamically changing subgraphs guided by the hand-crafted PSEs. In this way, our method can efficiently generate PSEs while maintaining the expressiveness. Finally, we propose an invariant learning module to learn graph representations that generalize to unknown test graphs through tailored objective functions. We also provide theoretical justifications for our proposed method in maintaining the sharpness of the attention function. We conduct extensive experiments on both synthetic and real-world graph benchmarks for the graph classification task. The results show that GOODFormer consistently and significantly outperforms various state-of-the-art baselines on unseen OOD testing graphs. Ablation studies and hyper-parameter analysis further demonstrate the designs of our proposed method.
@@ 3.1. Entropy-guided Invariant Subgraph Disentangler
[…]
@@
(5) $\boldsymbol{Z}_{\mathrm{MPNN},{\tilde{c}}}=\mathrm{MPNN}(\boldsymbol{Z},\boldsymbol{A}_{\tilde{c}}),\boldsymbol{Z}_{\mathrm{MPNN},{\tilde{s}}}=\mathrm{MPNN}(\boldsymbol{Z},\boldsymbol{A}_{\tilde{s}}),$
@@
(6) $\boldsymbol{Z}_{\tilde{c}}=\mathrm{MLP}(\boldsymbol{Z}_{\mathrm{Attn},{\tilde{c}}}+\boldsymbol{Z}_{\mathrm{MPNN},{\tilde{c}}}),$
$\boldsymbol{Z}_{\tilde{s}}=\mathrm{MLP}(\boldsymbol{Z}_{\mathrm{Attn},{\tilde{s}}}+\boldsymbol{Z}_{\mathrm{MPNN},{\tilde{s}}}),$
@@
[…]
More formally, we can prove the following theorem: {restatable}theoremsoftmaxdisperse Suppose $\boldsymbol{\mathrm{G}}_{S}\rightarrow\boldsymbol{\mathrm{G}}_{C}$ does not exist, $\mathcal{L}(\cdot)$ is a loss function and there exists one and only one non-trival subgraph $\boldsymbol{\mathrm{G}}_{C}$. Under these conditions, any model comprising Eq. (3) cannot guarantee the following intervention-based invariant learning principle: (1) minimizes $\mathbb{E}_{G_{s}}[\mathcal{L}(h(\boldsymbol{\mathrm{G}}),\boldsymbol{\mathrm{Y}}|do(\boldsymbol{\mathrm{G}}_{S}=G_{s}))]$, and simultaneously (2) minimizes $\mathrm{Var}_{G_{s}}(\{\mathcal{L}(h(\boldsymbol{\mathrm{G}}),\boldsymbol{\mathrm{Y}}|do(\boldsymbol{\mathrm{G}}_{S}=G_{s}))\})$, where $do(\boldsymbol{\mathrm{G}}_{S}=G_{s})$ denotes the do-calculus operation that removes every link from the parents to the variable $\boldsymbol{\mathrm{G}}_{S}$ and fixes $\boldsymbol{\mathrm{G}}_{S}$ to the specific value $G_{s}$. The proof is provided in Appendix C.2. The theorem shows that due to the limitations of the plain attention function, the disentangler cannot guarantee satisfying the principle upheld by the oracle $f_{Y}$ (given $\boldsymbol{\mathrm{G}}_{C}$), which restricts the OOD generalization capability. This limitation hampers the OOD generalization capability of our proposed disentangler under the invariant learning principle, which has shown effective within GNN architectures.
@@
where $t$ is the temperature. It is easy to see that the plain softmax function in Eq. (3) is equal to $t=1$. Using this function, we can reduce the temperature to decrease the entropy while maintaining the relative rank of the attention scores between pairs of nodes. Formally, we have the following theorem: {restatable}theoremtemperature Consider a vector $\boldsymbol{e}\in\mathbb{R}^{n}$, where $n>1$, comprising $n$ logits in the temperature-based softmax function Eq. (9) with $t>0$. Let $H=-\sum_{i}p_{i}\log p_{i},p_{i}=\mathrm{Softmax}_{t}(e_{i})$ denote the Shannon entropy of the temperature-based softmax function. As $t$ decreases, $H$ monotonically decreases. theorem The proof is provided in Appendix C.2. The theorem shows that we can adjust the entropy by changing the temperature. Therefore, during the testing stage, we optimize the temperature $t_{G}$ for each graph $G$. To prevent the model from being overly confident due to an excessively low temperature, we calculate the average entropy $\overline{H}$ during the training phase to guide the test-time training. The objective is to minimize the absolute difference between $\overline{H}$ and the entropy on test data, i.e.,:
@@
Combining the entropy-guided regularizer and temperature-based test time training, our disentangle can effectively distinguish invariant and variant subgraphs for subsequent components.
@@ 3.2. Evolving Subgraph Positional and Structural Encoder
(11) $\boldsymbol{Z}_{\mathrm{PSE},\tilde{c}}=h_{\mathrm{PSE}}(G_{\tilde{c}}^{\prime}),\boldsymbol{Z}_{\mathrm{PSE},\tilde{s}}=h_{\mathrm{PSE}}(G_{\tilde{s}}^{\prime}),$
where $G_{\tilde{c}}^{\prime}=(\boldsymbol{X},\mathcal{E}_{\tilde{c}})$ and $G_{\tilde{s}}^{\prime}=(\boldsymbol{X},\mathcal{E}_{\tilde{s}})$ represents the invariant and variant subgraphs with random node features, respectively. The edge sets $\mathcal{E}_{\tilde{c}}$ and $\mathcal{E}_{\tilde{s}}$ are kept the same as $\boldsymbol{A}_{\tilde{c}}$ and $\boldsymbol{A}_{\tilde{s}}$, respectively. To guide the encoder maximally learn positional and structural information, we also design a predictor $w_{\mathrm{PSE}}(\cdot)$ to recover the hand-crafted position encodings as follows:
(12) $\mathcal{L}_{\mathrm{PSE}}=\mathbb{E}_{G_{\tilde{c}}}[\lvert z_{\mathrm{PSE}}-w_{\mathrm{PSE}}(\boldsymbol{Z}_{\mathrm{PSE},\tilde{c}})\rvert]+\mathbb{E}_{G_{\tilde{s}}}[\lvert z_{\mathrm{PSE}}-w_{\mathrm{PSE}}(\boldsymbol{Z}_{\mathrm{PSE},\tilde{s}})\rvert],$
@@
(13) $\hat{y}_{\tilde{c}}=w_{\tilde{c}}(\mathrm{Pooling}(\mathrm{GT}^{\prime}([\boldsymbol{Z}_{\tilde{c}},\boldsymbol{Z}_{\mathrm{PSE},\tilde{c}}]))),$
$\hat{y}_{\tilde{s}}=w_{\tilde{s}}(\mathrm{Pooling}(\mathrm{GT}^{\prime}([\boldsymbol{Z}_{\tilde{s}},\boldsymbol{Z}_{\mathrm{PSE},\tilde{s}}]))),$
where $\mathrm{GT}^{\prime}(\cdot)$ represent hybrid MPNN and Transformer layers, $\mathrm{Pooling(\cdot)}$ is a global pooling function, $w_{\tilde{c}}(\cdot)$ and $w_{\tilde{s}}(\cdot)$ are two classifiers. Following (Wu et al., 2022a), during training, we formulate the overall prediction as
(14) $\hat{y}=\hat{y}_{\tilde{c}}\odot\sigma(\hat{y}_{\tilde{s}}),$
which mimics the intervention $do(S=\tilde{s})$ in causal theory. During the inference phase, we use $\hat{y}_{\tilde{c}}$ as the final prediction, which shields the influence of variant subgraph $G_{\tilde{s}}$.
@@ 3.3. Invariant GT Learning Module
Besides, we also have a normal classification loss function for the classifiers $w_{\tilde{s}}(\cdot)$ as:
@@ 13   Update model parameters using back propagation ;
We present the pseudo-code of GOODFormer in Algorithm 1. Additionally, the time complexity of our GOODFormer is $O(\lvert\mathcal{E}\rvert d+\lvert\mathcal{V}\rvert^{2}d^{2})$, where $\lvert\mathcal{V}\rvert$ and $\lvert\mathcal{E}\rvert$ denote the number of nodes and edges, respectively, and $d$ is the dimensionality of the representations. Specifically, we employ a combination of MPNN and Transformer layers to implement GOODFormer, resulting in a complexity of $O(\lvert\mathcal{E}\rvert d+\lvert\mathcal{V}\rvert^{2}d^{2})$. The time complexity during test-time training, which involves iterative computation of the softmax function, is $O(\lvert\mathcal{V}\rvert^{2}T)$, where $T$ denotes the number of iterations. Since $T$ is a small constant, the overall time complexity remains $O(\lvert\mathcal{E}\rvert d+\lvert\mathcal{V}\rvert^{2}d^{2})$. In comparison, the complexity of MPNNs is $O(\lvert\mathcal{E}\rvert d+\lvert\mathcal{V}\rvert d^{2})$ and the time complexity of representative GTs is $O(\lvert\mathcal{V}\rvert^{2}d^{2})$. Therefore, the complexity of GOODFormer is on the same scale as that of the existing GT methods. We also evaluate the computational overhead introduced by GOODFormer’s architecture and auxiliary objectives in Appendix E.2.
@@ 4.1.1. Datasets
We employ a synthetic dataset (GOOD-Motif) and four real-world datasets (GOOD-HIV, DrugOOD, GOOD-SST2, GOOD-Twitter) for the graph classification task (Gui et al., 2022). For the synthetic dataset, we use both the base and size splits of GOOD-Motif. For molecular property prediction tasks, we apply the scaffold and size splits of GOOD-HIV, and the assay split of DrugOOD to evaluate the performance of our method across various shifts. Additionally, GOOD-SST2 and GOOD-Twitter are graph-based natural language processing datasets, characterized by variations in sentence lengths. We report accuracy for GOOD-Motif, GOOD-SST2 and GOOD-Twitter, and ROC-AUC scores for GOOD-HIV and DrugOOD.
@@ 4.1.2. Compared Methods
We compare our GOODFormer with representative state-of-the-art methods from three groups. The first group is general invariant learning methods, including standard ERM, IRM (Arjovsky et al., 2019), and VREx (Krueger et al., 2021). The second group is graph OOD methods including DIR (Wu et al., 2022a), GIL (Li et al., 2022c), GSAT (Miao et al., 2022), CIGA (Chen et al., 2022b) and iMoLD (Zhuang et al., 2023). The third group is graph Transformer methods including Graphormer (Ying et al., 2021), GraphGPS (Rampášek et al., 2022), and EXPHORMER (Shirzad et al., 2023).
@@ 4.1.3. Optimization and Hyper-parameters
: For $\mathrm{GT}^{M}(\cdot)$ employing the GraphGPS layer (Rampášek et al., 2022), along with MPNN $h_{\mathrm{PSE}}(\cdot)$ and $\mathrm{GT}^{\prime}(\cdot)$, the numer of the respective layers is 2, 3 and 1, respectively. For all GTs, the dimensionality of both graph-level and node-level representations is $d=128$, whereas for GNN-based baselines, $d=300$. The global pooling function employed is mean pooling. The hand-crafted PSE $z_{\mathrm{PSE}}$ is Laplacian eigenvectors encoding (LapPE). The hyper-parameter $\lambda$ in Eq. (15) is chosen from $\{10^{2},10^{1},10^{0},10^{-1}\}$. The hyper-parameter $\alpha_{E}$ in Eq. (17) is chosen from $\{10^{-1},10^{-2}\}$. The hyper-parameter $\alpha_{\mathrm{PSE}}$ and $\alpha_{S}$ in Eq. (17) are set at $10^{-2}$ and $10^{0}$, repectively. We report the mean results and standard deviations of three runs. More details on the datasets, baselines, optimization and hyper-parameters are in Appendix D.

### v2

@@ Abstract.
journalyear: 2026copyright: ccconference: Proceedings of the 32nd ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1; August 09–13, 2026; Jeju Island, Republic of Koreabooktitle: Proceedings of the 32nd ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1 (KDD ’26), August 09–13, 2026, Jeju Island, Republic of Koreadoi: 10.1145/3770854.3780247isbn: 979-8-4007-2258-5/2026/08submissionid: 558ccs: Mathematics of computing Graph algorithmsccs: Computing methodologies Neural networks
Resource Availability:
The source code of this paper has been made publicly available at https://doi.org/10.5281/zenodo.18058301 and can also be accessed via the project repository: https://github.com/Lowy999/GOODFormer.
@@ 1. Introduction
To address these challenges, we propose GOODFormer, Graph Out-Of-Distribution Generalized Transformer . Our proposed method handles the distribution shifts of graphs by capturing invariant subgraphs and learning OOD generalized graph representations while resolving the aforementioned two challenges. Specifically, as shown in Figure 1, our proposed method achieves this goal by jointly optimizing three mutually reinforcing modules. First, we design a GT-based entropy-guided invariant subgraph disentangler to separate invariant and variant subgraphs using attention coefficients. To maintain the sharpness of the attention function, we propose entropy-guided sharp attention, comprising an entropy-guided regularizer and a lightweight temperature-based test-time training method. Next, we introduce an evolving subgraph positional and structural encoder to learn PSEs of dynamically changing subgraphs guided by the hand-crafted PSEs. In this way, our method can efficiently generate PSEs while maintaining the expressiveness. Finally, we propose an invariant learning module to learn graph representations that generalize to unknown test graphs through tailored objective functions. We also provide theoretical justifications for our proposed method in maintaining the sharpness of the attention function. We conduct extensive experiments on both synthetic and real-world graph benchmarks for the graph classification task. The results show that GOODFormer consistently and significantly outperforms various state-of-the-art baselines on unseen OOD testing graphs. Ablation studies and hyper-parameter analysis further demonstrate the designs of our proposed method.
@@ 3.1. Entropy-guided Invariant Subgraph Disentangler
where $\boldsymbol{Z}_{\mathrm{Attn},{\tilde{c}}}$ and $\boldsymbol{Z}_{\mathrm{Attn},{\tilde{s}}}$ are attention outputs for the invariant and variant subgraph, respectively, $\boldsymbol{Z}\in\mathbb{R}^{\lvert\mathcal{V}\rvert\times d}$ are $d$-dimensional intermediate node representations calculated by graph Transformer layers $\mathrm{GT}^{M}(\cdot)$ and $\boldsymbol{E}$ are intermediate attention scores calculated as $\boldsymbol{E}=\frac{\boldsymbol{Z}\boldsymbol{W}_{Q}(\boldsymbol{Z}\boldsymbol{W}_{K})^{\top}}{\sqrt{d_{K}}}$, $\boldsymbol{W}_{Q}$, $\boldsymbol{W}_{K}$ and $\boldsymbol{W}_{V}$ are three projection matrices, and $d_{K}$ is the hidden dimension. In short, these two attentions can distinguish invariant and variant subgraphs due to the complementarity design, i.e., if $(i,j)\in G_{\tilde{c}}$, then $\boldsymbol{E}_{(i,j)}>0$, and if $(i,j)\in G_{\tilde{s}}$, then $\boldsymbol{E}_{(i,j)}<0$, where $i$ and $j$ represent the indices of the nodes. The $\mathrm{softmax}$ function is used to highlight the node relationships $(i,j)$ that belong to $G_{\tilde{c}}$, while minimizing the scores of those node relationships $(i,j)$ that belong to $G_{\tilde{s}}$ as much as possible, ideally approaching zero. Compared to other invariant subgraphs extraction methods such as Top-K pooling (Zhang et al., 2018), softmax offers more flexibility and is gradient-friendly since it avoids any discrete operation. Besides, softmax is more adaptive at distinguishing invariant and variant subgraphs compared to the sigmoid function, which exhibits very small gradients when input values approach 0 or 1.
@@
[…]
@@
[…]
@@
where $\mathrm{Softmax}(\boldsymbol{E})_{i,j}$ and $\mathrm{Softmax}(\boldsymbol{-E})_{i,j}$ corresponds to the attention score of node relationship $(i,j)$ in $G_{c}$ and $G_{s}$ in Eq. (3), respectively, and $\epsilon$ is a small positive number that ensures the disentangler can correctly distinguish invariant and variant subgraphs.
More formally, we can prove the following theorem: {restatable}theoremsoftmaxdisperse Suppose $\boldsymbol{\mathrm{G}}_{S}\rightarrow\boldsymbol{\mathrm{G}}_{C}$ does not exist, $\mathcal{L}(\cdot)$ is a loss function and there exists one and only one non-trival subgraph $\boldsymbol{\mathrm{G}}_{C}$. Under these conditions, any model comprising Eq. (3) cannot guarantee the following intervention-based invariant learning principle: (1) minimizes $\mathbb{E}_{G_{s}}[\mathcal{L}(h(\boldsymbol{\mathrm{G}}),\boldsymbol{\mathrm{Y}}|do(\boldsymbol{\mathrm{G}}_{S}=G_{s}))]$, and simultaneously (2) minimizes $\mathrm{Var}_{G_{s}}(\{\mathcal{L}(h(\boldsymbol{\mathrm{G}}),\boldsymbol{\mathrm{Y}}|do(\boldsymbol{\mathrm{G}}_{S}=G_{s}))\})$, where $do(\boldsymbol{\mathrm{G}}_{S}=G_{s})$ denotes the do-calculus operation that removes every link from the parents to the variable $\boldsymbol{\mathrm{G}}_{S}$ and fixes $\boldsymbol{\mathrm{G}}_{S}$ to the specific value $G_{s}$. The proof is provided in Appendix D.2. The theorem shows that due to the limitations of the plain attention function, the disentangler cannot guarantee satisfying the principle upheld by the oracle $f_{Y}$ (given $\boldsymbol{\mathrm{G}}_{C}$), which restricts the OOD generalization capability. This limitation hampers the OOD generalization capability of our proposed disentangler under the invariant learning principle, which has shown effective within GNN architectures.
@@
where $t$ is the temperature. It is easy to see that the plain softmax function in Eq. (3) is equal to $t=1$. Using this function, we can reduce the temperature to decrease the entropy while maintaining the relative rank of the attention scores between pairs of nodes. Formally, we have the following theorem: {restatable}theoremtemperature Consider a vector $\boldsymbol{e}\in\mathbb{R}^{n}$, where $n>1$, comprising $n$ logits in the temperature-based softmax function Eq. (9) with $t>0$. Let $H=-\sum_{i}p_{i}\log p_{i},p_{i}=\mathrm{Softmax}_{t}(e_{i})$ denote the Shannon entropy of the temperature-based softmax function. As $t$ decreases, $H$ monotonically decreases. The proof is provided in Appendix D.2. The theorem shows that we can adjust the entropy by changing the temperature. Therefore, during the testing stage, we optimize the temperature $t_{G}$ for each graph $G$. To prevent the model from being overly confident due to an excessively low temperature, we calculate the average entropy $\overline{H}$ during the training phase to guide the test-time training. The objective is to minimize the absolute difference between $\overline{H}$ and the entropy on test data, i.e.,:
@@
Combining the entropy-guided regularizer and temperature-based test time training, our disentangler can effectively distinguish invariant and variant subgraphs for subsequent components.
@@ 3.2. Evolving Subgraph Positional and Structural Encoder
[…]
where $\lvert\cdot\rvert$ denotes the $\ell_{1}$ loss and $\boldsymbol{z}_{\mathrm{PSE}}$ is the hand-crafted position encoding of the original unmasked graph. This loss encourages the encoder to utilize the positional and structural information from the partially subgraph to approximate the PSE of the original graph. Note that $\boldsymbol{z}_{\mathrm{PSE}}$ can contain information from invariant subgraphs and lead to shortcuts. To prevent the introduction of spurious correlations into the disentangler, we have severed the back-propagation from the encoder to the disentangler.
@@
[…]
$\hat{y}_{\tilde{c}}$= $),\displaystyle w_{\tilde{c}}(\mathrm{Pooling}(\boldsymbol{Z}^{\prime}_{\tilde{c}}$)),
$\hat{y}_{\tilde{s}}$= $),\displaystyle w_{\tilde{s}}(\mathrm{Pooling}(\boldsymbol{Z}^{\prime}_{\tilde{s}}$)),
[…]
@@ 3.3. Invariant GT Learning Module
[…]
@@ 13   Update model parameters using back propagation ;
We present the pseudo-code of GOODFormer in Algorithm 1. Additionally, the time complexity of our GOODFormer is $O(\lvert\mathcal{E}\rvert d+\lvert\mathcal{V}\rvert^{2}d^{2})$, where $\lvert\mathcal{V}\rvert$ and $\lvert\mathcal{E}\rvert$ denote the number of nodes and edges, respectively, and $d$ is the dimensionality of the representations. Specifically, we employ a combination of MPNN and Transformer layers to implement GOODFormer, resulting in a complexity of $O(\lvert\mathcal{E}\rvert d+\lvert\mathcal{V}\rvert^{2}d^{2})$. The time complexity during test-time training, which involves iterative computation of the softmax function, is $O(\lvert\mathcal{V}\rvert^{2}T)$, where $T$ denotes the number of iterations. Since $T$ is a small constant, the overall time complexity remains $O(\lvert\mathcal{E}\rvert d+\lvert\mathcal{V}\rvert^{2}d^{2})$. In comparison, the complexity of MPNNs is $O(\lvert\mathcal{E}\rvert d+\lvert\mathcal{V}\rvert d^{2})$ and the time complexity of representative GTs is $O(\lvert\mathcal{V}\rvert^{2}d^{2})$. Therefore, the complexity of GOODFormer is on the same scale as that of the existing GT methods. We also evaluate the computational overhead introduced by GOODFormer’s architecture and auxiliary objectives in Appendix F.2.
@@ 4.1.1. Datasets
We employ a synthetic dataset (GOOD-Motif) and four real-world datasets (GOOD-HIV, DrugOOD, GOOD-SST2, and GOOD-Twitter) for the graph classification task (Gui et al., 2022). For the synthetic dataset, we use both the base and size splits of GOOD-Motif. For molecular property prediction tasks, we apply the scaffold and size splits of GOOD-HIV, and the assay split of DrugOOD to evaluate the performance of our method across various shifts. Additionally, GOOD-SST2 and GOOD-Twitter are graph-based natural language processing datasets, characterized by variations in sentence lengths. We report accuracy for GOOD-Motif, GOOD-SST2 and GOOD-Twitter, and ROC-AUC scores for GOOD-HIV and DrugOOD. The dataset statistics are presented in Table 1. More details on the datasets are in Appendix E.1.
Table 1. Statistics of datasets.
Dataset Number of Graphs
Train Validation Test
GOOD-Motif-basis/size 18,000 3,000 3,000
GOOD-HIV-scaffold 24,682 4,113 4,108
GOOD-HIV-size 26,169 4,112 3,961
GOOD-SST2-length 24,744 17,206 17,490
GOOD-Twitter-length 2,590 1,785 1,457
DrugOOD-assay 34,179 19,028 19,032
@@ 4.1.2. Compared Methods
We compare our GOODFormer with representative state-of-the-art methods from three groups. The first group is general invariant learning methods, including standard ERM, IRM (Arjovsky et al., 2019), and VREx (Krueger et al., 2021). The second group is graph OOD methods including DIR (Wu et al., 2022a), GREA (Liu et al., 2022), GIL (Li et al., 2022c), GSAT (Miao et al., 2022), CIGA (Chen et al., 2022b), iMoLD (Zhuang et al., 2023), and CAL+ (Sui et al., 2024). The third group is graph Transformer methods including Graphormer (Ying et al., 2021), GraphGPS (Rampášek et al., 2022), and EXPHORMER (Shirzad et al., 2023). More details on the baselines are in Appendix E.2.
@@ 4.1.3. Optimization and Hyper-parameters
For $\mathrm{GT}^{M}(\cdot)$ employing the GraphGPS layer (Rampášek et al., 2022), along with MPNN $h_{\mathrm{PSE}}(\cdot)$ and $\mathrm{GT}^{\prime}(\cdot)$, the numer of the respective layers is 2, 3 and 1, respectively. For all GTs, the dimensionality of both graph-level and node-level representations is $d=128$, whereas for GNN-based baselines, $d=300$. The global pooling function employed is mean pooling. The hand-crafted PSE $z_{\mathrm{PSE}}$ is Laplacian eigenvectors encoding (LapPE). The hyper-parameter $\lambda$ in Eq. (15) is chosen from $\{10^{2},10^{1},10^{0},10^{-1}\}$. The hyper-parameter $\alpha_{E}$ in Eq. (17) is chosen from $\{10^{-1},10^{-2}\}$. The hyper-parameter $\alpha_{\mathrm{PSE}}$ and $\alpha_{S}$ in Eq. (17) are set at $10^{-2}$ and $10^{0}$, repectively. We report the mean results and standard deviations of three runs. For the number of epochs, we use 100 for our method and the GT baselines. An exception is the GOOD-Motif basis split, which is limited to 50 epochs to prevent overfitting. Other GNN-based baselines adhere to the GOOD benchmark (Gui et al., 2022) and their respective original papers. We adopt Adaptive Moment Estimation (Adam) for optimization. The activation function is ReLU (Agarap, 2018). The selected hyper-parameters are reported in Table 2. Besides, our hyper-parameters have concrete meanings and do not necessitate a resource-intensive tuning:
[…]
GOOD-Motif GOOD-HIV GOOD-SST2 GOOD-Twitter DrugOOD
basis size scaffold size length length assay
$\lambda$ 10 1 0.1 100 0.1 100 100
$\alpha_{E}$ 0.1 0.1 0.1 0.01 0.01 0.01 0.01
$\alpha_{\mathrm{PSE}}$ 0.01 0.01 0.01 0.01 0.01 0.01 0.01